Find the point of equilibrium for a system that has a demand equation of p= 49-0,000 3x and a supply equation of p= 33+ 0.00002x
Answer by Boreal(15235)

Set the two equal, and I am going to edit what I think the numbers are.

49-0.00003x=33+0.00002x
16=0.00005x
x=320000
Check
49-0.00003(320000)=39.4
33+0.00002(320000)=39.4
If the numbers are different, then I would repost them.